OpenStudy (anonymous):

George makes large cookies. These cookies are 18 inches in diameter. How much icing will he need to cover the top of each cookie? 18π in2 324π in2 81π in2 9π in2

OpenStudy (anonymous):

@Jamierox4ev3r

OpenStudy (ivettef365):

you need to find the area A = pi.r^2

OpenStudy (anonymous):

what r and what does r^2 mean

jimthompson5910 (jim_thompson5910):

r is the radius r^2 means "r squared"

OpenStudy (anonymous):

oh yea my Bad

jimthompson5910 (jim_thompson5910):

that's ok, glad you asked

OpenStudy (ivettef365):

do you know what is r equal to ?

OpenStudy (anonymous):

would it be 324

jimthompson5910 (jim_thompson5910):

r is NOT 18

jimthompson5910 (jim_thompson5910):

18 is the diameter (not the radius)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

9

OpenStudy (ivettef365):

diameter = 18, and r (radius) equal half of 18

jimthompson5910 (jim_thompson5910):

good, r = 9 now square both sides

OpenStudy (anonymous):

81

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

right

OpenStudy (lncognlto):

81π in2

OpenStudy (anonymous):

thanks'

OpenStudy (anonymous):

@Hero

hero (hero):

All you need to do is use the formula \(A = \pi r^2\) You have to convert radius to diameter. \(r = \frac{d}{2}\) so \[A = \pi \left(\frac{d}{2}\right)^2\] d = 18 Input d into the formula and simplify

OpenStudy (anonymous):

18/2=9 9^2=81

hero (hero):

Don't ignore pi

OpenStudy (anonymous):

81pi